Evaluate the double integral (9 - x^2 - y^2)^(1/2) where


R = {(x,y) : x^2 + y^2 < or = to 9 , x > or = to 0 }


1
Expert's answer
2021-10-22T15:29:23-0400

ANSWER :9π

EXPLANATION

To calculate the integral , we transform the Cartesian coordinates into polar coordinates: "x=r cos\u03b8, \u03c5=r sin\u03b8" . The region of integration is a semicircle of radius "r=3, \theta \in \left[ -\frac { \pi }{ 2 } ,\frac { \pi }{ 2 } \right]" . Since "x^2+y^2=9" , then

"\iint _{ R }^{ \ }{ \sqrt { 9-{ x }^{ 2 }-{ y }^{ 2 } } dxdy\ =\int _{ -\frac { \pi }{ 2 } }^{ \ \frac { \pi }{ 2 } }{ \int _{ 0 }^{ 3 }{ r } \sqrt { 9-{ r }^{ 2 } } d\theta dr\ } }="

"=\left( \int _{ -\frac { \pi }{ 2 } }^{ \ \frac { \pi }{ 2 } }{ d\theta } \right) \cdot \left( \int _{ 0 }^{ 3 }{ r } \sqrt { 9-{ r }^{ 2 } } \ dr \right) =" "-\frac { \pi }{ 2 } \int _{ 0 }^{ 3 }{ \ } \sqrt { 9-{ r }^{ 2 } } \ d\left( 9-{ r }^{ 2 } \right) =-\frac { \pi }{ 2 } \cdot \frac { 2 }{ 3 } \ \left[ F(3)-F(0) \right]" . Because"F(r)={ \left[ { \left( 9-{ r }^{ 2 } \right) }^{ \frac { 3 }{ 2 } } \right] }", so "F(3)-F(0)=-27" and "\iint _{ R }^{ \ }{ \sqrt { 9-{ x }^{ 2 }-{ y }^{ 2 } } dxdy\ =9\pi \quad \quad \quad \quad \quad \quad \quad \quad \quad \quad }"